1. A gas expands from 282mL to 823mL at constant temperature. Calculate the work done (in joules) by the gas if it expands.

  a.) against a vacuum

  b.) against a constant pressure


1
Expert's answer
2022-03-19T02:44:21-0400

Work done by the gas in isothermal expansion w = -p∆V

p = pressure, ∆V = final volume (V2) - initial volume (V1)

(a) When expand against vacuum then p = 0 atm

      W = -0 atm×(0.823 - 0.282) L = 0 L atm = 0 J


(b) Here, p = const atm 

       V2-V1 = (0.823-0.282) L = 0.541 L

1 L atm = 101.325 J

W = - const×0.541 L atm